# Fundamental Oscillations of Massive Boson Stars - II

### v1

@@ 4.3 Universal Relations
4.3.1 Empirical Fits
The $f$-mode frequency is known to scale with the square root of average density ($M/R^{3}$) of NS and was first proposed by [58, 59]. The relation is of the form
$f(\text{kHz})=a+b\sqrt{\frac{GM}{R^{3}}}~{}.$ (4.3)
Here, $f$ is the $f$-mode frequency, $M$ is the mass of the star, and $R$ is its radius. $a$ and $b$ are the fitting coefficients.
This relation holds for ordinary NSs [59, 60], for NSs with hyperons [61], and also for DM admixed NS [62] as shown recently using the neutron decay model for DM. This relation was also checked for BSs for select values of $\lambda$ and $m$ in [31] that form BSs with masses in the range of NSs. The fit was performed for each set individually. The coefficients reported considering all the sets were $a=-0.0195\pm 0.0008$ kHz and $b=62.997\pm 0.058$ kHz km. The relation was also explored in [33] for select BS cases and reported the coefficients $a=-0.01461$ kHz and $b=42.11$ kHz km. Also, to compare with the coefficients for other NS cases, $a$ is known to lie within the range $0.22-0.79$ kHz and $b$ within $32-48$ kHz km. As these parameters are different from NSs, Flores et al. [31] concluded that these new relations can be used to distinguish BSs from NSs. We note that $a$ is small but negative in the case of BSs, and $b$ is higher compared to the other NS cases.
We reproduce $f$ as a function for $\sqrt{GM/R^{3}}$ for these cases as reported in Fig.3 of [31] (not shown here). There is a slight uncertainty involved as the curves do not identically overlap. As we discussed in this work, all the quantities scale with $x$. Hence, we check $f^{\prime}$ as a function of $\sqrt{M^{\prime}/R^{\prime 3}}$. We thus expect a unique relation as these quantities do not depend on the model parameters. This is shown in Fig. 5(a). We perform a linear fit given by
$f^{\prime}=-0.0045+0.21\sqrt{\frac{M^{\prime}}{R^{\prime 3}}}~{}.$ (4.4)
The fit agrees within $2.6\%$. We refer to the fit coefficients in Eq. 4.4 as $a^{\prime}=-00045$ and $b^{\prime}=0.21$. Writing back in unprimed quantities as given in Eq. 4.3, we get $b=b^{\prime}=0.21=63.0$ kHz km. Thus, $b$ remains unchanged, and we obtain the same slope. This value is consistent with the one obtained in [31] but is higher compared to the one obtained in [33]. $a=a^{\prime}/(xM_{Pl})=5.67\times 10^{-7}/(x/\text{MeV}^{-2})$ kHz.
This equation can be used to understand the dependence on these quasi-universal relations as a function of model parameters and explain the little variation obtained in [31]. For NS mass range as considered in [31, 33], $x\approx\Delta x\approx\mathcal{O}(10^{-5})$ MeV-2. Differentiating $a=a(x)$ with respect to $x$, we get $a\approx\Delta a\approx\mathcal{O}(10^{-2})$ kHz, which is approximately the value and error found in the parameter $a$ in [31, 33]. They also report an error in $b$; however, from this analysis, we expect $b$ to be invariant. The residual variations could be attributed to numerical error. We can also say that a variation of $\mathcal{O}(0.1)$ or higher in $a$ is expected for $x\lesssim 10^{-6}$. For larger values of $x$, i.e., BSs with higher maximum masses, the coefficient approaches zero asymptotically.
Another empirical relation for the $f$-mode damping time exists as a function of compactness when it is scaled by mass and radius [59]. The fit for BSs is given by
$\frac{R^{4}}{M^{3}\tau}=0.111-0.386C~{}.$ (4.5)
where, $M$, $R$, $C$, and $\tau$ are mass, radius, compactness and damping time of $f$-mode oscillations.The coefficients are in close agreement with the recent fit provided by [33]$:\frac{R^{4}}{M^{3}\tau}=0.1105-0.3764C~{}.$ The relation is invariant under the change to primed quantities. Thus, we directly plot $R^{4}/M^{3}\tau$ as a function of $C$ in Fig. 5(b). This is truly a universal relation for massive BSs and is followed no matter what the underlying model parameters are. We show the linear fit given by Eqn. 4.5 with a blue curve. The fit error is within $1.5\%$. Since the quantities plotted are invariant, we can directly compare them to the corresponding band obtained for NSs [59, 60, 61]. This band is shown in grey, and we conclude that BSs have a different relation. This also can be used to distinguish them from NSs. The fit given in [31] is not shown here. Fig. 3 of  [31], the compactness goes beyond the value of 0.18. However, the compactness for BSs, in our case, only goes up to 0.16. Also, they use a different fitting relation. Hence, we do make a comparison here.
@@ 4.3.1 Empirical Fits
Figure 5: Universal relations connecting (a) $f$-mode frequency in primed coordinates to the average density of BS and (b) scaled damping time $(R^{\prime 4}/M^{\prime 3}\tau^{\prime})$ to compactness as suggested in [59]. The grey band is the uncertainty in NSs with and without hyperonic matter. The band spans the fit relations reported in  [59, 60, 61]. The blue curves are fits given in Eqs. 4.4 adn 4.5. The errors in BS fits are shown in lower panels.
@@
There are also other relations involving mass-scaled $f$-mode characteristics that are universal in nature. The complex eigenfrequency, when scaled with mass, shows a tight correlation with compactness [59, 61]. This fit is of the quadratic form as in Eq. 4.1 with $Y=M\text{Re}(\omega_{f})$, $X=C$ and $Y=M\text{Im}(\omega_{f})/C^{4}$, $X=C$. Here, $\text{Re}(\omega_{f})=2\pi f$ is the $f$-mode angular frequency, and $\text{Im}(\omega_{f})=1/\tau$ is the inverse of damping time. Another universal relation exists between the mass-scaled complex $f$-mode frequency and the dimensionless tidal deformability [63, 64, 65]. The fit is of the quartic form of Eqn. 4.1 with $Y=M\omega_{f}$ and $X=\ln{\Lambda}$.
We plot these the relations for $\text{Re}(M\omega_{f})$ in Fig. 6 and for $\text{Im}(M\omega_{f})$ in Fig. 7 respectively. Note that $M\omega_{f}=M^{\prime}\omega^{\prime}_{f}$. All these quantities are independent of the model parameters $x$ and, again, are truly universal for the case of massive BSs. These are exact relations with no spread. These can also be compared with the case of NSs. Ref. [61] also obtained such fits using a large set of nuclear and hyperonic EoS. These are plotted in green-dashed lines in all the figures. The relations with the tidal deformability are consistent with the case of NSs. The ones containing compactness, however, show deviation and follow different relations than NSs. The fits for BSs, as obtained in this work, are shown in blue. The fit coefficients are tabulated in Table. 3 We see the $f$-Love and the $f-C$ fits have an error of less that $0.5\%$ and $1\%$ respectively. For the case of $\tau$ relations, the errors are higher, with it being within $4\%$ for $\tau-C$ and as high as $50\%$ for $\tau$-Love.
$Y$ $X$ $a_{0}$ $a_{1}$ $a_{2}$ $b_{0}$ $b_{1}$ $b_{2}$ $b_{3}$ $b_{4}$
$\text{Re}(M\omega_{f})$ $\ln{\Lambda}$ 2.365 ×$\times 10^{-1}$3.753 ×$\times 10^{-2}$ 2.103 ×$\times 10^{-3}$4.204 ×$\times 10^{-5}$ 5.420×$\times 10^{-8}$
$\text{Re}(M\omega_{f})$ $C$0.0024 0.237 1.886 - - - - -
$\text{Im}(M\omega_{f})$ $\ln{\Lambda}$ 2.024 ×$\times 10^{-4}$5.189 ×$\times 10^{-5}$ 5.000 ×$\times 10^{-6}$2.141 ×$\times 10^{-7}$ 3.428×$\times 10^{-9}$
@@
$N$ $1/C$ 1.833 1.442 ×$\times 10^{-1}$3.901 ×$\times 10^{-3}$ 6.121 ×$\times 10^{-5}$3.879 ×$\times 10^{-7}$
@@
Table 3: Fitting coefficients for the $f$-Love-$C$ universal relations and $N-C$. The fit equations are given in Eqs. 4.1.
@@
Figure 6: Universal relations for mass-scaled $f$-mode frequency as a function of the (a) logarithm of the dimensionless tidal deformability and (b) compactness. The fits are shown in blue, whose fitting coefficients are provided in Table. 3. The corresponding percentage errors are shown in the bottom panels. The relations followed by NS and hyperonic NSs are shown in green [61]
Figure 7: Same as Fig. 6 but for the $f$-mode damping time
@@
The quantity $N=f\tau=f^{\prime}\tau^{\prime}$ given the number of oscillations that the BS undergoes before its amplitude damp by a factor of $e$. Again, this quantity does not scale with $x$. We plot it against $C$ in Fig. 8 (taken from [34]). We see that $N$ decreases with compactness, i.e., more compact configurations for any set of $(\lambda,m)$ undergo less number of oscillations before damping out. For $C\approx 0.02$, $N$ can be as high as 10,000. For the case of the most compact configuration, we get a minimum number of such oscillations given by $N_{min}=406$. All BSs oscillate at least $N_{min}$ times within the damping time. This also translates to a high quality-factor ($Q=\pi f\tau$) of GW waveform, which will be discussed later in Sec. 4.5. We show a quartic polynomial (Eqn. 4.1) fit in blue. The fit has a percent error below $2\%$. The fit coefficients are reported in Table. 3.
Figure 8: Number of cycles ($N=f\tau=N^{\prime}\tau^{\prime}$) completed by BSs within damping time. BSs complete a minimum of $N_{min}=406$ oscillations for the case of the highest compactness configuration. The analytical fit is shown in blue, whose fitting coefficients are provided in Table. 3. The corresponding percentage error is shown in the bottom panels
4.4 Observable parameter space
@@
In this section, we focus on the scalar DM parameter space. We discussed various constraints in Sec. 3, and the available parameter space is visually shown in Fig. 1. For this entire region, it would be interesting to obtain the information of $f$-mode characteristics.
@@
First, we show the same parameter space with the constraints in Fig. 9. For every set of parameters ($m$, $\lambda$), we can compute the maximum possible $f$-mode frequency ($f_{max}$), corresponding to the most compact configuration. We can see that $f_{max}$ is higher for larger DM mass ($m$) and for smaller values of self-interaction ($\lambda$). Thus, the trend is inverse to that of the maximum mass $M_{\max}$. This trend is clear from the scaling relation $f_{max}=0.21/(xM_{Pl})=(0.21/M_{Pl})(m^{2}/\sqrt{\lambda})$ The variation is over orders of magnitude; hence we show all the quantities in log scale.
Figure 9: The maximum $f$-mode frequency for the entire parameter space for BSs. The colour indicates the $\log_{10}(f_{max})$ in Hz. The constraints from Fig. 1 are used to mask the excluded region in red. Only the unmasked region is relevant to the analysis. The frequency sensitivity bands for LIGO (green) and two of the future detectors, LISA (blue) and NEMO (orange) are shown. The frequency sensitivity assumed for LISA is 0.1 mHz - 1Hz, [50-1000] Hz for LIGO, and [2-4] kHz for NEMO. The region to the right of each band is sensitive to each detector.
@@
From an observation point of view, it would be of great interest to know what region of this vast parameter space could actually be accessible in the near future. $f$-mode oscillations of BSs would result in the emission of GW of $f$-mode frequencies. Hence, for reference, we plot the frequency band sensitive to various current and future GW detectors. We show the bands for three detectors, LISA, LIGO, and NEMO, in Fig. 9 in blue, green, and orange, respectively. The frequency sensitivity assumed is [0.1 mHz, 1 Hz] for LISA, [50, 1000] Hz for LIGO, and [2, 4] kHz for NEMO. These bands can be used as iso-contours for $f_{max}$ in the plot. The bands for Cosmic Explorer (CE) and Einstein Telescope (ET) are not explicitly shown. Their sensitivity bands encompass those of the LIGO and NEMO and further push down in the lower frequency range of a few Hz. Thus, all the GW detectors together roughly cover the region bounded by the blue and orange bands.
It is important to note that since we are plotting the maximum $f$-mode frequency here, it does not mean that these are the only bands that can be probed. The region towards the right of these bands has $f_{max}$ larger than the sensitivity range, i.e., the frequency of a configuration with lower compactness for these parameters can still fall in the detection band. Hence, in principle, the entire parameter space falling to the right of the band for each detector can be probed by that detector. This idea will be made clearer later when we show the exact masses and compactness that these GW detectors can probe.
Since we know that the $f_{max}$ scales with a particular combination of ($m$, $\lambda$), which is given by $f_{max}\propto m^{2}/\sqrt{\lambda}=1/x$, it is more meaningful to focus on the parameter $x$ and obtain the range of $x$ that can be probed. In Fig. 10, we show this relation of $f_{max}$ on $x$ in a black line. All axes are in log scale. $f_{max}$ decreases with $x$. The bands for the three detectors, as discussed, are shown in the same colour. Note that the black line marks the upper limit on frequency. We show $x$ here in the range [$10^{-6}$, $10^{3}$] MeV-2. For lower values of $x$, the compactness of BS would have to be too low to be detectable. We will show this in the next plot. We see that at $x=10^{3}$ MeV-2, the black line is already below the frequency range of all detectors and is not observable. The point where the black line goes below this detection threshold is $x_{thresh}=2.66\times 10^{2}$ MeV2. This is shown by the rightmost black-dashed line in the figure. This is an absolute upper threshold on $x$, such that $x>x_{thresh}$ is not accessible to the currently planned GW detectors. This threshold is set by the one observing in the lowest frequency range, i.e., LISA. The Pulsar Timing Array (PTA) observes GW in nano-Hz frequencies but is restricted to stochastic GW. Hence, we have not considered it here.
@@
Figure 10: Maximum $f$-mode frequency (black) and minimum damping time (red) as a function of scaling parameter. The corresponding maximum mass possible is shown on the upper axis. The frequency bands for LIGO and future planned GW detectors LISA and NEMO are shown in great, blue, and orange bands, respectively. The frequency sensitivity assumed for LISA is 0.1 mHz - 1Hz for LISA, [50, 1000] Hz for LIGO, and [2, 4] kHz for NEMO. The vertical dashed lines indicate the upper limit on $x$ that can be probed by NEMO, LIGO, and LISA, respectively, as we go from left to right. For higher values of $x$, the maximum $f$-mode frequency falls below the detector sensitivity range. The three values are $1.33\times 10^{-5}$ MeV-2, $5.32\times 10^{-4}$ MeV-2, and $2.66\times 10^{2}$ MeV-2 for NEMO, LIGO, and LISA, respectively.
@@
Similar upper thresholds can be obtained for other detectors as well. For LIGO, the threshold is $x\leq 5.32\times 10^{-4}$ MeV-2 while for NEMO it is $x\leq 1.33\times 10^{-5}$ MeV-2. If we consider the proposed deci-HZ detectors [66, 67] that are expected to be sensitive in $[0.1,10]$ Hz range, we obtain $x\leq 0.266$ MeV-2. In principle, the region $x\leq 1.33\times 10^{-5}$ MeV-2 is accessible to all detectors.
@@
On the same plot, we show some more information. As seen before, we can also calculate the minimum value of the damping time for a given $x$ as $\tau_{min}=1900xM_{Pl}$. We show this by the red curve, and the axis is marked on the right side of the figure. Since damping time is not easily measurable, we focus on the frequency for the rest of the study. On the upper axis, we show the values of maximum mass corresponding to the values of $x$ given by $M_{max}=0.06xM_{Pl}^{3}$. This figure compares the maximum frequency with the GW detector bands and marks the maximum mass of BS that would be observable. The threshold on $x$ would change if we considered BSs of lower compactness. In general, we would like to see the actual ranges of mass and $x$ that can be probed. This is studied in detail in the following subsection.
@@
In Fig. 11(a), we show the mass range corresponding to various values of $x$ that can be probed by the same GW detectors. The black line marks the maximum possible mass for a given value of $x$ ($M=0.06xM_{Pl}^{3}$), which also corresponds to the maximum compactness case. The region above this is inaccessible as we enter the unstable branch. These masses are expected to collapse to a BH. The dashed line marks the mass of BSs having $C=0.02$. As $M^{\prime}(C=0.02)\approx 0.012$, this corresponds to $M=0.012xM_{Pl}^{3}$. As we solve for $f$-modes only for $C\gtrsim 0.02$, we only focus on the regions between these two lines. Using the $f^{\prime}-M^{\prime}$ relation (Eq. 4.2), we can use the scaling to write $f=f(x,M)$. Doing this, for each $(x,M)$ in this region, we can check if the $f$-mode frequency falls in any of the GW detectors bands. The region in $(x,M)$ falling in LISA, LIGO, and NEMO bands are shown in blue, green, and orange, respectively. Thus for each $x$, we get a range in mass that can be probed. We see that BSs roughly in the mass range of $10^{2}-10^{7}M_{\odot}$ are observable by LISA, $0.1-10M_{\odot}$ by LIGO and $0.01-1M_{\odot}$ by NEMO. Note that for each case, we can probe even lower masses if we consider BSs with compactness lower than 0.02.
@@
To get a more physical picture, in Fig. 11(b), we show the same information in the $M-C$ plane. We make use of Eq.14 of [34] for obtain $f=f(C,M)$. Thus for each $(C,M)$ we can compare the BS $f$-mode frequency with various detector bands. The compactness varies between 0.02 and 0.16, the maximum for BS. We conclude that a mass range lower than that corresponding to $C_{max}$ can be probed for low compactness configurations. In this place, we can make a direct comparison with NSs as they typically lie within 1-3 $M_{\odot}$ and have compactness in the range 0.1-0.3. BHs fall on the extreme right of this diagram at $C=0.5$.
@@
Figure 11: (a) Masses of BSs falling in various GW bands for a given set of model parameters $x$. (b) Mass and compactness of BSs that could be probed by LIGO and planned future detectors LISA and NEMO. The mass-compactness range for NSs is marked by the grey patch. The frequency sensitivity assumed for LISA is 0.1 mHz-1Hz for LISA, [50, 1000] Hz for LIGO, and [2, 4] kHz for NEMO. The NS mass-compactness range is marked by the silver patch.

### v3

@@ 4.3 Universal Relations
Compact stars are known to follow some empirically observed universal relations, which can be used to infer stellar properties from observations of stellar oscillations. Such inference of stellar properties using observation of pulsations comes under the study of asteroseismology. In this section, we discuss some known universal relations focusing on the $f$-mode characteristics. Other known relations for static stars, in particular the II-Love-$C$, are discussed in Appendix. C where we use the scaling derived in Appendix. A and Appendix. B to show the II-LoveC-C universal relations are exact for massive BSs.
4.3.1 Empirical Fits
The $f$-mode frequency is known to scale with the square root of average density ($M/R^{3}$) of NSs and was first proposed by [AnderssonKokkotas1996, AnderssonKokkotas1998]. The relation is of the form
$f(\text{kHz})=a+b\sqrt{\frac{GM}{R^{3}}}~.$ (4.3)
Here, $f$ is the $f$-mode frequency, $M$ is the mass of the star, and $R$ is its radius. $a$ and $b$ are the fitting coefficients. This relation was checked for BSs for select values of $\lambda$ and $m$ in [Flores2019] that form BSs with masses in the range of NSs and the coefficients reported were $a=-0.0195\pm 0.0008$ kHz and $b=62.997\pm 0.058$ kHz km. The relation was also explored in [Celato2025] for select BS cases and reported the coefficients a=0.01461a=-0.01461 kHz and $b=42.11$ kHz km.
We reproduced the results from Fig.3 of [Flores2019], connecting $f$ with average density (not shown here), which reported a spread as the curves did not identically overlap. We checked that upon scaling the parameters and plotting the scaled quantities, all curves coincide to result in a unique curve, independent of DM model parameters. This is shown in Fig. 4(a). We perform a linear fit to this curve given by
$f^{\prime}=-0.0045+0.21\sqrt{\frac{M^{\prime}}{R^{\prime 3}}}~.$ (4.4)
The fit agrees within $2.6\%$. We refer to the fit coefficients in Eq. 4.4 as $a^{\prime}=-0.0045$ and $b^{\prime}=0.21$. Writing back in unprimed quantities as given in Eq. 4.3, we get $b=b^{\prime}=0.21=63.0$ kHz km. This value is consistent with the one obtained in [Flores2019] but is higher compared to the one obtained in [Celato2025]. This could be due to the difference in the density range for which the fit was performed and the small number of parameters and configurations that were considered in these works. In this work, we have derived this equation in an EoS-independent way rather than choosing some select model parameters and fitting them.
$a=a^{\prime}/(xM_{\rm Pl})=5.67\times 10^{-7}/(x/\text{MeV}^{-2})$ kHz depends on $x$. This dependence on $x$ can explain the variation obtained in [Flores2019] in the slope parameter. They also report an error in $b$; however, from this analysis, we expect $b$ to remain the same for all $x$, since $b^{\prime}=b$. The residual variations could be attributed to numerical error.
Another empirical relation for the $f$-mode damping time exists as a function of compactness when it is scaled by mass and radius [AnderssonKokkotas1998]. The fit for a BS is given by
$\frac{R^{4}}{(GM)^{3}\tau}=0.111-0.386C~.$ (4.5)
@@ 4.3.1 Empirical Fits
Figure 4: Universal relations connecting (a) $f$-mode frequency in primed coordinates to the average density of BS and (b) scaled damping time $(R^{\prime 4}/M^{\prime 3}\tau^{\prime})$ to compactness as suggested in [AnderssonKokkotas1998]. The grey band is the uncertainty in NSs with and without hyperonic matter. The band spans the fit relations reported in [AnderssonKokkotas1998, Benhar2004, Pradhan2022]. The blue curves are fits given in Eqs. 4.4 and 4.5. The errors in BS fits are shown in the lower panels.
4.3.2 ff-LoveC-C
There are also other relations involving mass-scaled $f$-mode characteristics that are universal in nature. The complex eigenfrequency, when scaled with mass, shows a tight correlation with compactness [AnderssonKokkotas1998, Pradhan2022] and the dimensionless tidal deformability [Chan2014, SotaniKumar2021, Pradhan2023]. The fits to these are of the quadratic and quartic forms, respectively. Here, $\text{Re}(\omega_{f})=2\pi f$ is the $f$-mode angular frequency, and $\text{Im}(\omega_{f})=1/\tau$ is the inverse of damping time.
We plot these the relations for $\text{Re}(M\omega_{f})$ in Fig. 5 (a) and (b) and for $\text{Im}(M\omega_{f})$ (not shown). It can be checked that $M\omega_{f}/M_{\rm Pl}^{2}=M^{\prime}\omega^{\prime}_{f}$. Thus, $M^{\prime}\omega^{\prime}_{f}$, just like $\bar{\Lambda}$ and $C$ do not scale with $x$. Since all these plotted quantities are independent of the model parameters, these relations are again universal, making the curves unique, i.e., these are followed by all massive BSs in the strong interaction limit. Ref. [Pradhan2022] also obtained such fits using a large set of nuclear and hyperonic EoS. We show these fits for comparison. The relations with the tidal deformability are consistent with the case of NSs. The ones containing compactness, however, show deviation and follow different relations than NSs. We see the $f$-Love and the fCf-C fits have an error of less than $0.5\%$ and $1\%$ respectively. For the case of $\tau$ relations, the errors are higher, within $4\%$ for $\tau-C$ and as high as $50\%$ for $\tau$-Love.
$Y$ $X$ $a_{0}$ $a_{1}$ $a_{2}$ $b_{0}$ $b_{1}$ $b_{2}$ $b_{3}$ $b_{4}$
@@
$\text{Re}(M\omega_{f})$ $C$ -0.0024 0.237 1.886 - - - - -
[…]
@@
NN $1/C$ - - - 1.833 1.442 $\times 10^{-1}$ -3.901 $\times 10^{-3}$ 6.121 $\times 10^{-5}$ -3.879 $\times 10^{-7}$
@@
Table 3: Fitting coefficients for the $f$-Love-$C$ universal relations and NCN-C. The fit equations are given in Eqs. 4.1.
@@
Figure 5: Universal relations for mass-scaled $f$-mode frequency as a function of the (a) logarithm of the dimensionless tidal deformability and (b) compactness. The relations followed by NSs and hyperonic NSs are shown in green [Pradhan2022]. (c) Number of cycles ($N=f\tau=N^{\prime}\tau^{\prime}$) completed by BSs within the damping time. BSs complete a minimum of $N_{\rm min}=406$ oscillations for the case of the highest compactness configuration. The NCN-C curves for pure neutron stars [Pradhan2022] are shown in grey for comparison. The analytical fits to the numerical solutions are shown in blue, whose fitting coefficients are provided in Table. 3. The corresponding percentage errors are shown in the bottom panels.
@@
The quantity $N=f\tau=f^{\prime}\tau^{\prime}$ gives the number of oscillations that the BS undergoes before its amplitude damps by a factor of $e$. Again, this quantity does not scale with $x$. We plot it against $C$ in Fig. 5(c) (taken from [shirke_BS_fmodes_short]). We see that $N$ decreases with compactness, i.e., more compact configurations undergo fewer oscillations before damping out. For $C\approx 0.02$, $N$ can be as high as 10,000. For the case of the most compact configuration, we get a minimum number of such oscillations given by $N_{\rm min}=406$. We show a quartic polynomial (Eqn. 4.1) fit, with a fitting error below $2\%$. Since $f^{\prime}\tau^{\prime}=f\tau$, we compare them to the corresponding band obtained for NSs [Pradhan2022].
4.4 Observable DM parameter space
@@
In this section, we focus on the scalar DM parameter space. We discussed various constraints in Sec. 3, and the available parameter space is visually shown in Fig. 1. For this entire region, it would be interesting to obtain the information of $f$-mode characteristics and check which regions can be probed by various GW detectors.
@@
First, we show the same parameter space with the constraints in Fig. 6(a). For every set of parameters ($m$, $\lambda$), we can compute the maximum possible $f$-mode frequency ($f_{\rm max}$), corresponding to the most compact configuration. We can see that $f_{\rm max}$ is higher for larger DM mass ($m$) and for smaller values of self-interaction ($\lambda$). This trend is clear from the scaling relation $f_{\rm max}=0.21/(xM_{\rm Pl})=(0.21/M_{\rm Pl})(m^{2}/\sqrt{\lambda})$.
@@
From an observation point of view, it would be of great interest to know what region of this vast parameter space could actually be accessible in the near future. $f$-mode oscillations of BSs would result in the emission of GWs of $f$-mode frequencies. Hence, for reference, we plot the frequency band sensitive to various current and future GW detectors. We show the bands for three detectors, LISA, LIGO, and NEMO, in Fig. 6(a). The frequency range over which the detectors have high sensitivity for various detectors is taken to be [0.1 mHz, 1 Hz] for LISA, [50, 1000] Hz for LIGO, and [2, 4] kHz for NEMO. These bands are used as iso-contours for $f_{\rm max}$ in the plot. The bands for Cosmic Explorer (CE) and Einstein Telescope (ET) are not explicitly shown as their sensitivity bands roughly encompass those of LIGO and NEMO. Thus, all the GW detectors together roughly cover the region bounded by the blue and orange bands.
@@
Figure 6: a) The maximum $f$-mode frequency for the entire parameter space for BSs. The colourbar indicates the $\log_{10}(f_{\rm max})$ in Hz. The constraints from Fig. 1 are used to mask the excluded region in red. Only the unmasked region is relevant to the analysis. b) Maximum $f$-mode frequency (black) and minimum damping time (red) as a function of scaling parameter. The corresponding maximum mass possible is shown on the upper axis. The vertical dashed lines indicate the upper limit on $x$ that can be probed by NEMO, LIGO, and LISA, respectively, as we go from left to right. The frequency bands for LIGO and future planned GW detectors, LISA and NEMO, are overlaid. See text for more details.
It is important to note that since we are plotting the maximum $f$-mode frequency here, it does not mean that these are the only bands that can be probed. The region towards the right of these bands has $f_{\rm max}$ larger than the sensitivity range, i.e., the frequency of a configuration with lower compactness for these parameters can still fall in the detection band. Hence, in principle, the entire parameter space falling to the right of the band for each detector can be probed by that detector.
Since we know that $f_{\rm max}$ scales with a particular combination of ($m$, $\lambda$), given by $f_{\rm max}\propto m^{2}/\sqrt{\lambda}=1/x$, we focus on the parameter $x$, as we would like to infer microscopic DM parameters from future GWs from $f$-modes, and obtain the range of $x$ that can be probed. In Fig. 6(b), we show this relation of $f_{\rm max}$ on $x$ in a black line. Note that the black line marks the upper limit on frequency. We see that at $x=10^{3}$ MeV-2, the black line is already below the frequency range of all detectors and is not observable. The point where the black line goes below this detection threshold is $x_{\rm thresh}=2.66\times 10^{2}$ MeV2. This is shown by the rightmost black-dashed line in the figure. This is an absolute upper threshold on $x$, such that $x>x_{\rm thresh}$ is inaccessible to the currently planned GW detectors. This threshold is set by the one observing in the lowest frequency range, i.e., LISA. The Pulsar Timing Array (PTA) observes GWs in nano-Hz frequencies but is restricted to stochastic GWs. Hence, we have not considered it here.
@@
Similar upper thresholds can be obtained for other detectors as well. For LIGO, the threshold is $x\leq 5.32\times 10^{-4}$ MeV-2 while for NEMO it is $x\leq 1.33\times 10^{-5}$ MeV-2. If we consider the proposed deci-Hz detectors [decigo2011, tiango2020] that are expected to be sensitive in $[0.1,10]$ Hz range, we obtain $x\leq 0.266$ MeV-2. In principle, the region $x\leq 1.33\times 10^{-5}$ MeV-2 is accessible to all detectors.
@@
On the same plot, we also show the minimum value of the damping time for a given $x$ as $\tau_{\rm min}=1900xM_{\rm Pl}$ in red. Since the damping time is not easily measurable, we focus on frequency in this study. On the upper axis, we show the values of maximum mass corresponding to the values of $x$ given by $M_{\rm max}=0.06xM_{\rm Pl}^{3}$. The range of $x$ sensitive to a particular detector would change if we considered BSs of lower compactness. In general, we would like to see the actual ranges of mass and $x$ that can be probed.
@@
In Fig. 7(a), we show the mass range corresponding to various values of $x$ that can be probed by the same GW detectors. The black line marks the maximum possible mass for a given value of $x$ ($M=0.06xM_{\rm Pl}^{3}$), which also corresponds to the maximum compactness case. The region above this is inaccessible as we enter the unstable branch. These masses are expected to collapse into a BH. The dashed line marks the mass of BSs having a fixed $C=0.02$. Since we solve for $f$-modes (as mentioned in Sec. 2) only for $C\gtrsim 0.02$, we only focus on the regions between these two lines. Using the $f^{\prime}-M^{\prime}$ relation (Eq. 4.2), we can use the scaling to write $f=f(x,M)$. Doing this, for each $(x,M)$ in this region, we can check if the $f$-mode frequency falls in any of the GW detector bands. Such regions in $(x,M)$ falling in LISA, LIGO, and NEMO bands are shown in blue, green, and orange, respectively. We see that BSs roughly in the mass range of $10^{2}-10^{7}M_{\odot}$ are observable by LISA, $0.1-10M_{\odot}$ by LIGO and $0.01-1M_{\odot}$ by NEMO. Note that for each case, we can probe even lower masses if we consider BSs with compactness lower than 0.02.
@@
To get a more physical picture, in Fig. 7(b), we show the same information in the MCM-C plane. We make use of Eq.14 of [shirke_BS_fmodes_short] to obtain $f=f(C,M)$ and repeat the process. We conclude that a mass range lower than that corresponding to $C_{\rm max}$ can be probed for low compactness configurations. We can make a direct comparison with NSs as they typically lie within 1-3 $M_{\odot}$ and have compactness roughly in the range 0.1-0.3. BHs fall on the extreme right of this diagram at $C=0.5$.
@@
Figure 7: (a) Masses of BSs falling in various GW bands for a given set of model parameters $x$. (b) Mass and compactness of BSs that could be probed by LIGO and planned future detectors, LISA and NEMO. The mass-compactness range for NSs is marked by the silver patch.